Brazilia Bus Tours has incurred the following bus maintenance costs during the recent tourist season. (The real is Brazil’s national monetary unit. On the day this exercise was written, the real was equivalent in value to .5092 U.S. dollar.)


  Month Miles Traveled
by Tour Buses
Maintenance
Cost
  November 14,850 20,150 real    
  December 17,300 18,800          
  January 19,750 18,950         
  February 23,900 19,400          
  March 31,400 20,150          
  April 13,400 17,900         


Required:
1.

Use the high-low method to estimate the variable cost per tour mile traveled and the fixed cost per month. (Round "Variable cost per tour mile" to 3 decimal places.)

      

Variable cost per tour mile real
Fixed cost per month real
2.

Develop a formula to express the cost behavior exhibited by the company’s maintenance cost. (Round "Variable cost per tour mile" to 3 decimal places.)

      

Variable cost per tour mile real
Fixed cost per month real
3.

Predict the level of maintenance cost that would be incurred during a month when 43,000 tour miles are driven.

     

Maintenance cost real
Answer #1

1) Variable cost per tour mile = (20150-17900)/(31400-13400) = 0.125 per real

Fixed cost = 20150-(31400*.125) = $16225

2) Y = a+bx

Y = 16225+0.125X

3) Maintenance cost = (43000*.125)+16225 = $21600